(ii)The consumer faces the budget line P1X1 + P2X2= M where P1 and P2 are price for good 1 and 2 and X1and X2 are quantity demanded for good 1 and good 2 respectively , Mis consumer income. If the price of good 1 doubles, the price of good 2 becomes 5 times larger, and income becomes 3 times larger, write down an equation for the new budget line in terms of the original prices and income.



1
Expert's answer
2021-06-28T17:15:02-0400

Solution:

The budget line is P1X1 + P2X2= M where P1 and P2 are prices of commodities 1&2 and X1and X2 are the quantities demanded for goods 1&2 respectively.


When Price of good 1 doubles the equation becomes: 2( P1X1) + P2X2= M

When Price of good 2 increases five fold the equation becomes: 2( P1X1) + 5(P2X2)= M

When Income of consumer increases three fold the equation become: 2( P1X1) + 5(P2X2)=3(M)


Answer:

Thus, the new budget constraint becomes as follows:

2P1X1+5P2X2=3M


where P1 and P2 are initial prices of the initial goods 1 and 2 and initial money income would be equal to money income M.